### v1

# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Navigation

Abstract
Multi-agent reinforcement learning (MARL) has recently excelled in solving challenging cooperative and competitive multi-agent problems in various environments with, mostly, few agents and full observability. Moreover, a range of crucial robotics-related tasks, such as multi-robot navigation and obstacle avoidance, that have been conventionally approached with the classical non-learnable methods (e.g., heuristic search) is currently suggested to be solved by the learning-based or hybrid methods. Still, in this domain, it is hard, not to say impossible, to conduct a fair comparison between classical, learning-based, and hybrid approaches due to the lack of a unified framework that supports both learning and evaluation. To this end, we introduce POGEMA, a set of comprehensive tools that includes a fast environment for learning, a generator of problem instances, the collection of pre-defined ones, a visualization toolkit, and a benchmarking tool that allows automated evaluation. We introduce and specify an evaluation protocol defining a range of domain-related metrics computed on the basics of the primary evaluation indicators (such as success rate and path length), allowing a fair multi-fold comparison. The results of such a comparison, which involves a variety of state-of-the-art MARL, search-based, and hybrid methods, are presented.
(a) task
(b) observation
(c) results
Figure 1: (a) Example of the multi-robot navigation problem considered in POGEMA: each robot must reach its goal, denoted by a flag of the same color. (b) Observation tensor of the red agent. (c) Results of the evaluation of several MARL, hybrid, and search-based solvers on the proposed POGEMA benchmark.
1 Introduction
Multi-agent reinforcement learning (MARL) has gained an increasing attention recently and significant progress in this field has been achieved [6, 32, 62]. MARL methods have been demonstrated to generate well-performing agents’ policies in strategic games [2, 63], sport simulators [49, 66], multi-component robot control [59], city traffic control [20], and autonomous driving [70]. Currently, several ways to formulate and solve MARL problems exist, based on what information is available to the agents and what type of communication is allowed in the environment [67]. Due to the increased interest in robotic applications, decentralized cooperative learning with minimizing communication between agents has recently attracted a specific attention [45, 68]. Decentralized learning naturally suits the partial observability of the environment in which the robots usually operate. Reducing the information transmitted through the communication channels between the agents increases their degree of autonomy.
The main challenges in solving MARL problems are the non-stationarity of the multi-agent environment, the need to explicitly predict the behavior of the other agents to implement cooperative behavior, high dimensionality of the action space, which grows exponentially with the number of agents, and the sample inefficiency of existing approaches. The existing MARL including model-based and hybrid learnable methods [12, 26] exhibit faster and more stable learning in SMAC-type environments [13] with vector observations and full observability. Currently, the best results are shown by the discrete explicit world models, that use Monte Carlo tree search for planning with various heuristics to reduce the search space [18, 26].
However, in numerous practically inspired applications, like in mobile robot navigation, agents’ observations are typically high-dimensional (e.g. stacked occupancy grid matrices or image-based observations as compared to 32-dim vectors in SMAC [13]) and only partially describe the state of the environment, including the other agents [17, 14]. This makes the problem specifically challenging, especially in the environments where a large number of agents are involved. For example, it is not uncommon in robotics to consider settings where up to hundreds of agents are acting (moving) simultaneously in the shared workspace as opposed to 2–10 agents in conventional MARL environments such as SMAC [13] or Google Research Football [21]. Learning to act in such crowded, observation-rich and partially-observable environments is a notable challenge to existing MARL methods.
Conventionally, the problem of multi-robot cooperative navigation (which is very important due to its applications in modern automated warehouses and fulfillment centers [11]) is framed as a search problem over a discretized search space, composed of robots-locations tuples. All robots are assumed to be confined to a graph, typically – a 4-connected grid [38], and at each time step a robot can either move following a graph’s edge or stay at the current vertex. This problem setting is known as (Classical) Multi-agent Pathfinding problem [50]. Even in such simplified setting (discretized space, discretized time, uniform-duration actions etc.) obtaining a set of individual plans (one for each robot) that are mutually-conflict-free (i.e. no vertex or edge is occupied by disctinct agents at the same time step) and minimize a common objective such as, for example, the arrival time of the last agent (known as the makespan in the literature) is NP-Hard [65]. Moreover if the underlying graph is directed even obtaining a valid solution is HP-Hard as well [31].
To this end the focus of the multi-agent pathfinding community is recently being shifted towards exploring of how state-of-the-art machine learning techniques, especially reinforcement learning and imitation learning, can be leveraged to increase the efficiency of traditional solvers. Methods like [48, 46, 47, 29, 61, 42, 58, 27, 10] are all hybrid solvers that rely on both widespread search-based techniques and learnable components as well. They all are developed using different frameworks, environments and datasets and are evaluated accordingly, i.e. in the absence of the unifying evaluation framework, consisting of the (automated) evaluation tool, protocol (that defines common performance indicators) and the dataset of the problem instances. Moreover, currently most of the pure MARL methods, i.e. the ones that do not involve search-based modules, such as QMIX [37], MAMBA [12], MAPPO [64] etc., are mostly not included in comparison. The main reason is that to train MARL policies a fast environment is needed, which is suited to cooperative multi-agent navigation.
To close the mentioned gaps we introduce POGEMA, a comprehensive set of tools that includes:
  • a fast and flexible environment for learning and planning supporting several variants of the multi-robot navigation problem,
  • a generator of problem instances for multi-task and generalization testing,
  • a visualization toolkit to create plots for debugging and performance information and to make high-quality animations,
  • a benchmarking tool that allows automated evaluation of both learnable, planning, and hybrid approaches.
Moreover, we introduce and specify an evaluation protocol defining a range of domain-related metrics computed on the basics of the primary evaluation indicators (such as success rate and path length), allowing a fair multi-fold comparison of learnable and classical methods. The results of such a comparison, which involves a range of the state-of-the-art MARL, search-based, and hybrid methods, are presented. The source code for POGEMA Benchmark is available at the following links [^1] https://github.com/AIRI-Institute/pogema-benchmark [^2] https://github.com/AIRI-Institute/pogema-toolbox [^3] https://github.com/AIRI-Institute/pogema. Additionally, the POGEMA environment and pogema-toolbox can be installed via PyPI using pip.
2 Related Work
Currently, a huge variety of MARL environments exists that are inspired by various practical applications and encompass a broad spectrum of nuances in problem formulations. Notably, they include a diverse array of computer games [41, 13, 39, 7, 51, 19, 5, 3, 21]. Additionally, they address complex social dilemmas [1] including public goods games, resource allocation problems [35], and multi-agent coordination challenges. Some are practically inspired, showcasing tasks such as competitive object tracking [34], infrastructure management and planning [23], and automated scheduling of trains [30]. Beyond these, the environments simulate intricate, interactive systems such as traffic management and autonomous vehicle coordination [57], multi-agent control tasks [39, 36], and warehouse management [16]. Each scenario is designed to challenge and analyze the collaborative and competitive dynamics that emerge among agents in varied and complex contexts. We summarize the most wide-spread MARL environments in Table 1. A detailed description of each column is presented below.
Table 1: Comparison of different multi-agent reinforcement learning environments
Environment Repository
Navigation
Partially observable
Python based
Hardware-agnostic setup
Performance >10K Steps/s
Procedural generation
Requires generalization
Evaluation protocols
Tests & CI
PyPi Listed
Scalability >1000 Agents
Induced behavior
Flatland [30] link Coop
GoBigger [69] link Mixed/Coop
Google Research Football [21] link Mixed
Griddly [4] link Mixed
Hide-and-Seek [3] link Comp
IMP-MARL [23] link Coop
Jumanji (XLA) [5] link Mixed
LBF [35] link Coop
MAMuJoCo [36] link Coop
MATE [34] link Coop
MeltingPot [1] link Mixed/Coop
Minecraft MALMO [19] link Mixed
MPE [28] link Mixed
MPE (XLA) [39] link Mixed
Multi-agent Brax (XLA) [39] link Coop
Multi-Car Racing [44] link Comp
Neural MMO [51] link Comp
Nocturne [57] link Mixed
Overcooked [7] link Coop
Overcooked (XLA) [39] link Coop
RWARE [35] link Coop
SISL [16] link Coop
SMAC [41] link Mixed/Coop
SMAC v2 [13] link Mixed/Coop
SMAX (XLA) [39] link Mixed/Coop
POGEMA (ours) link Mixed
Navigation
Navigation tasks arise in almost all multi-agent environments (e.g. unit navigation in SMAC or robotic warehouse management in RWARE), however only a handful of environments specifically focus on challenging navigation problems: Flatland, Nocturne, RWARE, and POGEMA.
Partially observable
Partial observability is an intrinsic feature of a generic multi-agent problem, meaning that an individual agent does not have access to the full state of the environment but rather is able to observe it only locally (e.g. an agent is able to determine the locations of the other agents and/or static obstacles only in its vicinity). Most of the considered environments are partially observable, with the exception of Overcooked.
Python based
Most of the current multi-agent environments are implemented in Python, however, some of them include bindings of the other program languages or external dependencies that might complicate their usage. Pure Python implementations generally ensure ease of modification and customization, allowing researchers to readily adapt and extend the environments.
Hardware-agnostic setup
means that the environment doesn’t require setup for any specific type of training or inferencing hardware.
Performance >10K Steps/s
Training and evaluating multi-agent reinforcement learning agents often requires making billions of steps (transitions) in the environment. Thus, it is crucial that each transition is computed efficiently. In general, performing more than 10K steps per second is a good indicator of the environment’s efficiency. While XLA versions can provide high performance by vectorizing the environment on GPU, they require modern hardware setups, which can be a barrier for some researchers. In contrast, fast environments like POGEMA or RWARE can achieve high performance without such stringent hardware requirements, making them more accessible and easier to integrate into a variety of research projects.
Procedural generation
To improve the ability of RL agents to solve problem instances that are not the same that were used for training (the so-called ability to generalize) procedural generation of the problem instances is commonly used. I.e. the environment does not rely on a predefined set of training instances but rather procedurally generates them to prevent overfitting. As highlighted in the Procgen paper [9], this approach ensures that agents develop robust strategies capable of adapting to novel and diverse situations. Moreover, in multi-agent settings, agents must be able to handle and adapt to a variety of unforeseen agent behaviors and strategies, ensuring robustness and flexibility in dynamic environments [1].
Evaluation protocols
means that the environment features a comprehensive evaluation API, including computation of distinct performance indicators and visualization tools. These capabilities allow precise performance measurement and deeper insights into RL agents’ behavior, going beyond just reward curves, which can often hide agents exploiting the reward system rather than genuinely solving the tasks.
Tests and CI
means that the environment is set up for development with continuous integration and is covered with tests, which are essential for collaborative open-source development.
PyPI listed
means that the environment library is listed on PyPI [^4] https://pypi.org, making it easy to install and integrate into your projects with a simple pip install command.
Scalability to >1000 Agents
stands for the environment’s ability to handle more than 1000 simultaneously acting agents, ensuring robust performance and flexibility for large-scale multi-agent simulations.
Induced behaviour
Multi-agent behaviour can be induced by modifying the reward function. Competitive (Comp) behaviour occurs when all agents’ rewards sum to a fixed constant at each time step. As an example, minimax games, where rewards sum up to zero, are competitive games. Cooperative/collaborative (Coop) behaviour happens when the agents share a reward function – e.g. when both agents have the same goal in the environment and get rewarded only by its completion. Mixed behaviour (Mixed) does not constrain the reward function in any way and can be described as a combination of cooperative and competitive behaviours. One of the most famous examples of these games is iterated prisoner’s dilemma.
As we aim to create a lightweight and easy-to-configure multi-agent environment for reinforcement learning and pathfinding tasks, we consider the following factors essential. First and foremost, our environment is fully compatible with the native Python API: we target pure Python builds independent of hardware-specific software with a minimal number of external dependencies. Moreover, we underline the importance of constant extension and flexibility of the environment. Thus, we prioritize testing and continuous integration as cornerstones of the environment, as well as trouble-free modification of the transition dynamics. Secondly, we highlight that our environment targets generalization and may utilize procedural generation. Last but not least, we target high computational throughput (i.e., the number of environment steps per second) and robustness to an extremely large number of agents (i.e., the environment remains performant under high loads).
There are many environments inducing various types of multi-agent behaviors via different reward structures. Unfortunately, many of them require extensive Python support and rely on APIs of different programming languages (e.g., Lua, C++) for lower latency or depend on hardware-specific libraries such as XLA. Furthermore, many environments do not support generalization and lack procedural generation, especially in multi-agent cases. Additionally, customization of certain environments might be considered an issue without reverse engineering them. That’s why we emphasize the superiority of the proposed benchmark.
Despite the diversity of available environments, most research papers tend to utilize only a selected few. Among these, the most popular are the StarCraft Multi-agent Challenge (SMAC), Multi-agent MuJoCo (MAMuJoCo), and Google Research Football (GRF), with SMAC being the most prevalent in top conference papers. The popularity of these environments is likely due to their effective contextualization of algorithms. For instance, to demonstrate the advantages of a method, it is crucial to test it within a well-known environment.
The evaluation protocols in these environments typically feature learning curves that highlight the performance of each algorithm under specific scenarios. For SMAC, these scenarios involve games against predefined bots with specific units on both sides. In MAMuJoCo, the standard tasks involve agents controlling different sets of joints, while in GRF, the scenarios are games against predefined policies from Football Academy scenarios. Proper evaluation of MARL approaches is a serious concern. For SMAC, it’s highlighted in the paper [15], which proposes a unified evaluation protocol for this benchmark. This protocol includes default evaluation parameters, performance metrics, uncertainty quantification, and a results reporting scheme.
The variability of results across different studies underscores the importance of a well-defined evaluation protocol, which should be developed alongside the presentation of the environment. In our study, we provide not only the environment but also the evaluation protocol, popular MARL baselines, and modern learnable MAPF approaches to better position our benchmark within the context.
3 POGEMA
POGEMA, which comes from Partially-Observable Grid Environment for Multible Agents, is an umbrella name for a collection of versatile and flexible tools aimed at developing, debugging and evaluating different methods and policies tailored to solve several types of multi-agent navigation tasks.
3.1 POGEMA Environment
POGEMA [^5] https://pypi.org/project/pogema environment is a core of POGEMA suite. It implements the basic mechanics of agents’ interaction with the world. The environment can be installed using the Python Package Index (PyPI). The environemnt is open-sourced and available at github [^6] https://github.com/AIRI-Institute/pogema under MIT license. POGEMA provides integration with existing RL frameworks: PettingZoo [54], PyMARL [40], and Gymnasium [56].
Basic mechanics
The workspace where the agents navigate is represented as a grid composed of blocked and free cells. Only the free cells are available for navigation. At each timestep each agent individually and independently (in accordance with a policy) picks an action and then these actions are performed simultaneously. POGEMA implements collision shielding mechanism, i.e. if an agent picks an action that leads to an obstacle (or out-of-the-map) than it stays put, the same applies for two or more agents that wish to occupy the same cell. POGEMA also has an option when one of the agents deciding to move to the common cell does it, while the others stay where they were. The episode ends when the predefined timestep, episode length, is reached. The episode can also end before this timestep if certain conditions are met, i.e. all agents reach their goal locations if MAPF problem (see below) is considered.
Problem settings
POGEMA supports two generic types of multi-agent navigation problems. In the first variant, dubbed MAPF (from Multi-agent Pathfinding), each agent is provided with the unique goal location and has to reach it avoiding collisions with the other agents and static obstacles. For MAPF problem setting POGEMA supports both stay-at-target behavior (when the episode successfully ends only if all the agents are at their targets) and disappear-at-target (when the agent is removed from the environment after it first reaches its goal). The second variant is a lifelong version of multi-agent navigation and is dubbed accordingly – LMAPF. Here each agent upon reaching a goal is immediately assigned another one (not known to the agent beforehand). Thus the agents are constantly moving trough in the environment until episode ends.
Observation
At each timestep each agent in POGEMA receives an individual ego-centric observation represented as a tensor – see Fig. 1. The latter is composed of the following $(2R+1)\times(2R+1)$ binary matrices, where $R$ is the observation radius set by the user:
  1. 1.
    Static Obstacles – 0 means the free cell, 1 – static obstacle
  2. 2.
    Other Agents – 0 means no agent in the cell, 1 – the other agent occupies the cell
  3. 3.
    Targets – projection of the (current) goal location of the agent to the boundary of its field-of-view
The suggested observation, which is, indeed, minimalist and simplistic, can be modified by the user using wrapper mechanisms. For example, it is not uncommon in the MAPF literature to augment the observation with additional matrices encoding the agent’s path-to-goal (constructed by some global pathfinding routine) [46] or other variants of global guidance [29].
Reward
POGEMA features the most intuitive and basic reward structure for learning. I.e. an agent is rewarded with $+1$ if it reaches the goal and receives 00 otherwise. For MARL policies that leverage centralized training a shared reward is supported, i.e. $r_{t}=goals/agents$ where $goals$ is the number of goals reached by the agents at timestep $t$ and $agents$ is the number of agents. Indeed, the user can specify its own reward using wrappers.
Performance indicators
The following performance indicators are considered basic and are tracked in each episode. For MAPF they are: Sum-of-costs (SoC) and makespan. The former is the sum of time steps (across all agents) consumed by the agents to reach their respective goals, the latter is the maximum over those times. The lower those indicators are the more effectively the agents are solving MAPF tasks. For LMAPF the primary tracked indicator is the throughput which is the ratio of the number of the accomplished goals (by all agents) to the episode length. The higher – the better.
3.2 POGEMA Toolbox
The POGEMA Toolbox is a comprehensive framework designed to facilitate the testing of learning-based approaches within the POGEMA environment. This toolbox offers a unified interface that enables the seamless execution of any learnable MAPF algorithm in POGEMA. Firstly, the toolbox provides robust management tools for custom maps, allowing users to register and utilize these maps effectively within POGEMA. Secondly, it enables the concurrent execution of multiple testing instances across various algorithms in a distributed manner, leveraging Dask [^7] https://github.com/dask/dask for scalable processing. The results from these instances are then aggregated for analysis. Lastly, the toolbox includes visualization capabilities, offering a convenient method to graphically represent aggregated results through detailed plots. This functionality enhances the interpretability of outcomes, facilitating a deeper understanding of algorithm performance.
POGEMA Toolbox offers a dedicated tool for map generation, allowing the creation of three distinct types of maps: random, mazes and warehouse maps. All generators facilitates map creation using adjustable parameters such as width, height, and obstacle density. Additionally, maze generator includes specific parameters for mazes such as the number of wall components and the length of walls. The maze generator was implemented based on the generator provided in  [10]. POGEMA Toolbox [^8] https://pypi.org/project/pogema-toolbox can be installed using PyPI, and licenced under Apache License 2.0.
3.3 Baselines
POGEMA integrates a variety of MARL, hybrid and planning-based algorithms with the environment. These algorithms, recently presented, demonstrate state-of-the-art performance in their respective fields. Table 2 highlights the differences between these approaches. Some, such as LaCAM and RHCR, are centralized search-based planners. Other approaches, such as SCRIMP and DCC, while decentralized, still require communication between agents to resolve potential collisions. Learnable modern approaches for LifeLong MAPF that do not utilize communication include Follower [46], MATS-LP [47], and Switchers [48] (Assistant Switcher, Learnable Switcher). All these approaches utilize independent PPO [43] as the training method.
The following modern MARL algorithms are included as baselines: MAMBA [12], QPLEX [60], IQL [53], VDN [52], and QMIX [37]. For environment preprocessing, we used the preprocessing scheme provided in the Follower approach, enhancing it with the anonymous targets of other agents’ local observations. We utilized the official implementation of MAMBA, as provided by its authors [^9] https://github.com/jbr-ai-labs/mamba, and employed PyMARL2 framework [^10] https://github.com/hijkzzz/pymarl2 for establishing MARL baselines.
Table 2: This table provides an overview of various baseline approaches supported by POGEMA and their features in the context of decentralized multi-agent pathfinding.
Algorithm Decentralized Partial Observability Fully Integrated into POGEMA Supports MAPF Supports LifeLong MAPF No Global Obstacles Map No Communication Parameter Sharing Decentralized Learning Model-Based No Imitation Learning
MAMBA [12]
QPLEX [60]
IQL [53]
VDN [52]
QMIX [37]
Follower [46]
MATS-LP [47]
Switcher [48]
SCRIMP [61]
DCC [29]
LaCAM [33] - - - - -
RHCR [25] - - - - -
4 Evaluation Protocol
4.1 Dataset
We include the maps of the following types in our evaluation dataset (with the intuition that different maps topologies are necessary for proper assessment):
  • Mazes – maps that encouter prolonged corridors with 1-cell width that require high level of cooperation between the agent to accomplish the mission. These maps are procedurally generated.
  • Random – one of the most commonly used type of maps, as they are easy to generate and allow to avoid overfitting to some special structure of the map. POGEMA ontains an integrated random maps generator, that allows to control the density of the obstacles.
  • Warehouses – this type of maps are usually used in the papers related to LifeLong MAPF. While there is no narrow passages, high density of the agents might significantly reduce the overall throughput, especially when agents are badly distributed along the map. These maps are also can be procedurally generated.
  • MovingAI – a set of maps from the existing benchmark widely used in MAPF community. The contained maps have different sizes and structures. It can be used to show how the approach deals with single-agent pathfinding and also deals with the maps that have out-of-distribution structure.
  • MovingAI-tiles – a modified MovingAI set of maps. Due to the large size of the original maps, it’s hard to get high density of the agents on them. To get more crowded maps, we slice the original maps on 16 pieces with $64\times 64$ size.
  • Puzzles – a set of small hand-crafted maps that contains some difficult patterns that mandate the cooperation between that agents.
(a) maze
(b) random
(c) warehouse
Figure 2: Examples of map generators presented in POGEMA.
Start and goal locations are generated via random generators. They are generated with fixed seeds, thus can be reproduced. It’s guaranteed, that each agent has its own goal location and the path to it from its start location exists. Examples of the maps are presented in Figure 2.
4.2 Metrics
The existing works related to solving MAPF problems evaluates the performance by two major criteria – success rate and the primary performance indicators mentioned above: sum-of-costs, makespan, throughput. These are directly obtainable from POGEMA. While these metrics allow to evaluate the algorithms at some particular instance, it’s might be difficult to get a high-level conclusion about the performance of the algorithms. Thus, we want to introduce several high-level metrics that covers multiple different aspects:
Performance – how well the algorithm works compared to other approaches. To compute this metric we run the approaches on a set of maps similar to the ones, used during training, and compare the obtained results with the best ones.
$Performance_{MAPF}=\begin{cases}SoC_{best}/SoC\\ 0\text{ if not solved}\end{cases}$ (1)
$Performance_{LMAPF}=throughput/throughput_{best}$ (2)
Out-of-Distribution – how well the algorithm works on out-of-distribution maps. This metric is computed in the same way as Performance, with the only difference that the approaches are evaluated on a set of maps, that were not used during training phase and have different structure of obstacles. For this purpose we utilize maps from MovingAI-tiles set of maps.
$Out\_of\_Distribution_{MAPF}=\begin{cases}SoC_{best}/SoC\\ 0\text{ if not solved}\end{cases}$ (3)
$Out\_of\_Distributionn_{LMAPF}=throughput/throughput_{best}$ (4)
Scalability – how well the algorithm scales to large number of agents. To evaluate how well the algorithm scales to large number of agents, we run it on a large warehouse map with increasing number of agents and compute the ratio between runtimes with various number of agents.
$Scalability=\frac{runtime(agents_{1})/runtime(agents_{2})}{|agents_{1}|/|agents_{2}|}$ (5)
Cooperation – how well the algorithm is able to resolve complex situations. To evaluate this metric we run the algorithm on Puzzles set of maps and compare the obtained results with best solutions that were obtained by classical MAPF/LMAPF solvers.
$Cooperation_{MAPF}=\begin{cases}SoC_{best}/SoC\\ 0\text{ if not solved}\end{cases}$ (6)
$Cooperation_{LMAPF}=throughput/throughput_{best}$ (7)
Congestion – how well the algorithm distributes the agents along the map and reduces redundant waits, collisions, etc. To evaluate this metric we compute the average density of the agents presented in the observations of each agent and compare it to the overall density of the agents on the map.
$Congestion=\frac{\sum_{i\in agents}agents\_density(obs_{i})/agents\_density(map)}{|agents|}$ (8)
Pathfinding – how well the algorithm works in case of presence of a single agent on a large map. This metric is tailored to determine the ability of the approach to effectively lead agents to their goal locations. For this purpose we run the approaches on large city maps from MovingAI benchmark sets. The obtained solution cost (in fact - length of the path) should be optimal.
$Pathfinding=\begin{cases}1\text{ if path is optimal}\\ 0\text{ otherwise}\end{cases}$ (9)
4.3 Experimental Results
We have evaluated all the supported baseline algorithms (12 in total) on both MAPF and LMAPF setups on all 6 datasets. The results of this evaluation are presented in Fig.3. The details about number of maps, number of agents, seeds, etc. are given in the supplementary material (as well as details on how these results can be reproduced).
Figure 3: Evaluation of baselines available in POGEMA on (a) MAPF (b) LMAPF instances.
In both setups, i.e. MAPF and LMAPF, the best results in terms of cooperation, out-of-distribution and performance metrics were obtained by centralized planners, i.e. LaCAM and RHCR respectively.
For MAPF tasks, LaCAM outperforms all other approaches on all metrics except congestion. It is hypothesized that in this approach, the even distribution of agents across the environment is not crucial due to its centralized nature, which efficiently resolves complex conflicts. Specialized learnable MAPF approaches, i.e., DCC and SCRIMP, take second place, showing close performance but with different specifics. DCC shows better results on out-of-distribution tasks and pathfinding tasks than SCRIMP, which is better at managing congestion. Surprisingly, the results of SCRIMP are inferior on pathfinding tasks, suggesting a problem with this approach in single-agent tasks that do not require communication, which can be an out-of-distribution setup for this algorithm. MARL algorithms such as QPLEX, VDN, and QMIX underperform in comparison with other approaches, exhibiting a significant gap in the results, which can be attributed to the absence of additional techniques used in hybrid approaches, despite incorporating preprocessing techniques from the Follower approach. This could suggest that the MARL community lacks large-scale approaches and benchmarks for them. Predictably, IQL shows the poorest performance, highlighting the importance of centralized training for multi-agent pathfinding (MAPF) tasks that require high levels of cooperation. The best results out of all MARL approaches are shown by MAMBA, which has non-zero results on the performance metric and better results in terms of the cooperation metric. Its results are still much worse than those obtained by hybrid methods, and it is not able to solve any instances in the out-of-distribution dataset.
For LMAPF tasks, the centralized approach, RHCR, dominates in cooperation, out-of-distribution, performance, and pathfinding metrics. However, it significantly lags behind other approaches in terms of congestion and scalability metrics. Out of the non-centralized approaches, the best results, depending on the metric, are shown by either MATS-LP or Follower. While MATS-LP is better in terms of cooperation and pathfinding, the latter is better in terms of congestion and scalability metrics. It’s also worth noting that MATS-LP requires much more runtime than other approaches as it runs MCTS for each of the agents at each step, which takes time (see Appendix F for more details). There are also two more hybrid approaches – ASwitcher and LSwitcher – which differ in the way they switch between planning-based and learning-based parts. One of the reasons for their mediocre results is a total lack of global information, i.e., the Switcher approach assumes that the agents have no information about the global map; thus, each of them needs to reconstruct the map based on their local observations. Again, MARL approaches underperform in these scenarios, with QMIX and QPLEX showing comparable results. QMIX performs better in cooperation and out-of-distribution metrics, while QPLEX excels in performance. As in the MAPF setup, among all the MARL approaches, the best results are demonstrated by MAMBA, which is slightly better than others in terms of cooperation and out-of-distribution metrics and much better in terms of the performance metric.
5 Conclusion and Limitations
This paper presents POGEMA – a powerful suite of tools tailored for creating, assessing, and comparing methods and policies in multi-agent navigation problems. POGEMA encompasses a fast learning environment and a comprehensive evaluation toolbox suitable for pure MARL, hybrid, and search-based solvers. It includes a wide array of methods as baselines. The evaluation protocol described, along with a rich set of metrics, assists in assessing the generalization and scalability of all approaches. Visualization tools enable qualitative examination of algorithm performance. Integration with the well-known MARL API and map sets facilitates the benchmark’s expansion. Existing limitations are two-fold. First, a conceptual limitation is that communication between the agents is not currently disentangled in POGEMA environment. Second, the technical limitations include the lack of JAX support and integration with other well-known GPU parallelization tools.
Appendix Contents
Appendix Sections Contents
Appendix A
Evaluation Setup Details
Appendix B
Results for MAPF Benchmark
Appendix C
Results for LifeLong MAPF Benchmark
Appendix D
Code examples for POGEMA
Appendix E
POGEMA Toolbox
Appendix F
Extended Related Work
Appendix G
Examples of Used Maps
Appendix H
MARL Training Setup
Appendix I
Resources and Statistics
Appendix J
Community Engagement and Framework Enhancements
Appendix A Evaluation Setup Details
POGEMA benchmark contains $6$ different sets of maps and all baseline approaches were evaluated on them either on MAPF or on LMAPF instances. Regardless the type of instances, number of maps, seeds and agents were the same. Table 3 contains all information about these numbers. In total, this corresponds to 3,376 episodes for each scenario type. Note that there is no MaxSteps (LMAPF) value for MovingAI set of maps. This set of maps was used only for pathfinding meta-metric, thus all approaches were evaluated only on MAPF instances with a single agent.
Table 3: Details about the instances on different sets of maps.
Agents Maps MapSize Seeds MaxSteps MaxSteps
(MAPF) (LMAPF)
Random [8, 16, 24, 32, 48, 64] 128 17$\times$17 - 21$\times$21 1 128 256
Mazes [8, 16, 24, 32, 48, 64] 128 17$\times$17 - 21$\times$21 1 128 256
Warehouse [32, 64, 96, 128, 160, 192] 1 33$\times$46 128 128 256
Puzzles [2, 3, 4] 16 5$\times$5 10 128 256
MovingAI [1] 8 256$\times$256 10 2048 -
MovingAI-tiles [64, 128, 192, 256] 128 64$\times$64 1 256 256
Appendix B Results for MAPF Benchmark
In this section, we present the extended results of the MAPF benchmark analysis, highlighting the performance, out-of-distribution handling, scalability, cooperation, congestion, and pathfinding capabilities of various approaches. The experiments were conducted on different map types and sizes, employing metrics such as SoC, CSR, and makespan to evaluate effectiveness. Detailed visual and tabular data illustrate how centralized and learnable approaches perform under various conditions.
B.1 Performance
The performance metrics were calculated using Mazes and Random maps of size close to $20\times 20$. The primary metrics here are SoC and CSR. The results of all the MAPF approaches over different numbers of agents are presented in Figure 4. The superior performance is shown by the centralized approach, LaCAM. The learnable approaches, DCC and SCRIMP, show comparable results. Interestingly, the former has a better SoC metric, despite the latter having better results on CSR. Among the MARL methods, better results are shown by MAMBA for both metrics. However, it narrowly lags behind the specialized approaches, DCC and SCRIMP.
Figure 4: Performance of MAPF approaches on Random and Mazes maps, based on CSR (higher is better) and SoC (lower is better) metrics. The shaded area indicates $95\%$ confidence intervals.
B.2 Out-of-Distribution
Out-of-Distribution metric was calculated on MovingAI-tiles dataset, that consists of pieces of cities maps with $64\times 64$ size. Due to much larger size compared to Mazes and Random maps, the amount of agents was also significantly increased. The results are presented in Figure 5. Here again centralized search-based planner, i.e. LaCAM, demonstrates the best results both in terms of CSR and SoC. Among hybrid methods, DCC completely outperforms the others (including SCRIMP) on the out-of-distribution dataset. MARL approaches are unable to solve any instance even with 64 agents.
Figure 5: Performance of MAPF approaches on MovingAI-tiles maps. These results were utilized to compute Out-of-Distribution metric. The shaded area indicates $95\%$ confidence intervals.
B.3 Scalability
Figure 6: Runtime in seconds for each algorithm. The plot is log-scaled.
The results of how well the algorithm scales with a large number of agents are shown in Figure 6. The experiments were conducted on a warehouse map. The plot is log-scaled. The best scalability is achieved with the centralized LaCAM approach, which is a high-performance approach. The worst results in both runtime and scalability are for SCRIMP, with results close to it for DCC. Despite an initially high runtime, the scalability of MAMBA is better than other approaches; however, this could be attributed to the high cost of GPU computation, which is due to the large number of parameters in the neural network and is the limiting factor of this approach.
B.4 Cooperation
Table 4: Comparison of algorithms’ cooperation on Puzzles set. $\pm$ shows confidence intervals $95\%$. Here, tan boxes highlight the best approach, and teal boxes highlight the best approach with a learnable part.
Algorithm CSR SoC
DCC $0.72$ p m 0.04 $96.33$ p m 9.97
IQL $0.27$ p m 0.04 $316.87$ p m 14.47
LaCAM $0.96$ p m 0.02 $36.29$ p m 7.26
MAMBA $0.39$ p m 0.04 $177.64$ p m 14.68
QMIX $0.53$ p m 0.05 $246.11$ p m 17.10
QPLEX $0.33$ p m 0.04 $250.12$ p m 14.80
SCRIMP $0.82$ p m 0.04 $104.31$ p m 13.73
VDN $0.11$ p m 0.03 $336.99$ p m 12.45
How well the algorithm is able to resolve complex situations on the Puzzles set is reflected in the results presented in Table 4. Surprisingly, the centralized approach LaCAM does not solve all the tasks, showing only a $0.96$ CSR. This highlights that this type of task is difficult even for centralized approaches, despite the small map size of $5\times 5$ and the low number of agents ($2-4$). SCRIMP outperformed DCC in CSR but again showed comparable results in SoC. Among MARL approaches, better cooperation is demonstrated by QMIX, outperforming QPLEX, VDN, IQL, and even MAMBA.
B.5 Congestion
Congestion is one of the meta-metrics that estimates how well the agents are distributed along the map. This metric indirectly influences the performance of the approach, as well-distributed agents in case of highly crowded instances allows to reduce the amount of collisions and redundant wait-actions. To compute this metric we utilize the results obtained on Warehouse map with the highest evaluated amount of agents - $192$. In Table 5, we present results for a range of agents from $64$ to $192$ in increments of $32$, not just for $192$ agents. In contrast to other metrics, that are computed as a ratio to the best obtained results, Congestion metric is computed as the ratio to average density of all agents on the map to average density of the agents in agents’ local observations.
Table 5: Average agents density by number of agents across algorithms for Warehouse map. Please note, only column with $192$ agents was utilized to compute Congestion metric.
Algorithm 64 Agents 96 Agents 128 Agents 160 Agents 192 Agents
DCC $0.094$ p m 0.001 $0.132$ p m 0.001 $0.170$ p m 0.001 $0.207$ p m 0.001 $0.241$ p m 0.001
IQL $0.076$ p m 0.001 $0.114$ p m 0.001 $0.163$ p m 0.002 $0.244$ p m 0.002 $0.299$ p m 0.002
LaCAM $0.082$ p m 0.001 $0.116$ p m 0.001 $0.149$ p m 0.001 $0.179$ p m 0.001 $0.207$ p m 0.001
MAMBA $0.101$ p m 0.001 $0.183$ p m 0.001 $0.266$ p m 0.002 $0.335$ p m 0.002 $0.389$ p m 0.002
QMIX $0.073$ p m 0.001 $0.103$ p m 0.001 $0.130$ p m 0.001 $0.154$ p m 0.001 $0.179$ p m 0.001
QPLEX $0.077$ p m 0.001 $0.113$ p m 0.001 $0.146$ p m 0.001 $0.175$ p m 0.001 $0.205$ p m 0.001
SCRIMP $0.074$ p m 0.001 $0.104$ p m 0.001 $0.127$ p m 0.001 $0.148$ p m 0.001 $0.173$ p m 0.001
VDN $0.071$ p m 0.001 $0.101$ p m 0.001 $0.130$ p m 0.001 $0.158$ p m 0.001 $0.188$ p m 0.001
B.6 Pathfinding
Table 6: Comparison of makespan (the lower is better) used for pathfinding metric. tan boxes highlight the best approach, and teal boxes highlight the best approach with a learnable part.
Algorithm Makespan
DCC $189.56$ p m 28.28
IQL $1825.95$ p m 137.11
LaCAM $179.82$ p m 20.21
MAMBA $416.45$ p m 139.34
QMIX $955.54$ p m 203.76
QPLEX $933.74$ p m 204.21
SCRIMP $1460.04$ p m 176.27
VDN $1733.20$ p m 158.91
To compute Pathfidning metric we run the approaches on the instances with a single agent. For this purpose we utilized large MovingAI mapf with $256\times 256$ size, the results are presented in Table 6. While this task seems easy, most of the hybrid and MARL approaches are not able to effectively solve them. Only LaCAM is able to find optimal paths in all the cases, as it utilizes precomputed costs to the goal location as a heuristic. Most of the evaluated hybrid and MARL approaches are also contain a sort of global guidance in one the channels of their observations. However, large maps with out-of-distribution structure, the absence of communication and other agents in local observations are able to lead to inconsistent behavior of the models that are not able to effectively choose the actions that lead the agent to its goal. Please note, SoC and makespan metrics in this case are equal, as there is only one agent in every instance.
Appendix C Results for LifeLong MAPF Benchmark
In this section, we present the extended results of the LifeLong MAPF benchmark analysis, highlighting performance, out-of-distribution handling, scalability, cooperation, congestion, and pathfinding.
C.1 Performance
Performance metric in LMAPF case is based on the ratio of throughput compared to the best obtained one. In contrast to SoC, throughput should be as high as possible. There is also no CSR metric, as there is no need for agents to be at their goal locations simultaneously. As well as in MAPF case, the best results are obtained by centralized search-based approach – RHCR. The best results among decentralized methods demonstrate Follower and MATS-LP, Following them, comparable results are shown by ASwitcher and LSwitcher. Between MARL methods the highest throughput on both Random and Mazes maps is obtained by MAMBA.
Figure 7: Performance results for LifeLong scenarios on the Mazes and Random maps.
C.2 Out-of-Distribution
The evaluation on out-of-distribution set of maps confirms the results obtained on Random and Mazes maps. The best results demonstrates RHCR. Next best results are obtained by Follower and MATS-LP, which are much closer to RHCR in this experiment. While MATS-LP outperforms Follower on the instances with 64, 128 and 192 agents, Follower is still better on the instances with 256 agents. Such relation is probably explained by the presence of dynamic edge-costs in Follower that allows to better distribute agents along the map and reduce congestion between them.
Table 7: Evaluation on Out-of-Distribution maps. tan boxes highlight the best approach according to the average throughput metric, and teal boxes highlight the best approach with a learnable component.
Algorithm 64 Agents 128 Agents 192 Agents 256 Agents
ASwitcher $1.26$ p m 0.08 $2.30$ p m 0.13 $3.14$ p m 0.17 $3.80$ p m 0.20
Follower $1.50$ p m 0.08 $2.82$ p m 0.13 $3.95$ p m 0.19 $4.81$ p m 0.22
IQL $0.26$ p m 0.02 $0.65$ p m 0.04 $0.78$ p m 0.05 $0.68$ p m 0.05
LSwitcher $1.23$ p m 0.07 $2.23$ p m 0.12 $3.06$ p m 0.17 $3.67$ p m 0.20
MAMBA $1.02$ p m 0.05 $1.42$ p m 0.08 $2.05$ p m 0.12 $2.46$ p m 0.17
MATS-LP $1.57$ p m 0.12 $2.98$ p m 0.20 $4.04$ p m 0.33 $4.69$ p m 0.39
QMIX $0.83$ p m 0.03 $1.55$ p m 0.06 $2.01$ p m 0.09 $2.27$ p m 0.11
QPLEX $0.79$ p m 0.03 $1.48$ p m 0.07 $1.79$ p m 0.10 $1.74$ p m 0.11
RHCR $1.57$ p m 0.08 $3.00$ p m 0.14 $4.22$ p m 0.23 $5.13$ p m 0.34
VDN $0.50$ p m 0.03 $0.91$ p m 0.05 $1.03$ p m 0.06 $0.96$ p m 0.06
C.3 Scalability
Figure 8: Runtime in seconds for each algorithm. Note that the plot is log-scaled.
Figure 8 contains log-scaled plot of average time spent by each of the algorithms to process an instance on Warehouse map with the corresponding amount of agents. Most of the approaches scales almost linearly, except RHCR. This centralized search-based method lacks of exponential grow, as it needs to find a collision-free solution for at least next few steps, rather than just to make a decision about next action for each of the agents. The worst runtime demonstrate MATS-LP, as it runs MCTS and simulates the behavior of the other observable agents. It’s still scales better than RHCR as it builds trees for each of the agents independently.
C.4 Cooperation
Table 8: Average throughput on Puzzles maps that were used to compute Cooperation metric.
Algorithm Average Throughput
ASwitcher $0.164$ p m 0.015
Follower $0.319$ p m 0.020
IQL $0.036$ p m 0.003
LSwitcher $0.206$ p m 0.013
MAMBA $0.133$ p m 0.014
MATS-LP $0.394$ p m 0.021
QMIX $0.117$ p m 0.010
QPLEX $0.051$ p m 0.006
RHCR $0.538$ p m 0.021
VDN $0.030$ p m 0.004
As well as for MAPF setting, cooperation metric is computed based on the results obtained on Puzzles dataset. Table 8 contains average throughput obtained by each of the evaluated approaches. Here again the best results are obtained by RHCR algorithm. In contrast to Random, Mazes and Warehouse sets of maps, where MATS-LP and Follower demonstrate close results, the ability to simulate the behavior of other agents, provided by MCTS in MATS-LP, allows to significantly outperform Follower on small Puzzles maps. The rest approaches demonstrate much worse results, especially IQL, QPLEX and VDN that have 10 times worse average throughput than RHCR.
C.5 Congestion
Table 9 contains average agents density presented in observations. As it was already mentioned in LMAPF:Out-of-Distribution section, Follower contains a mechanism that allows to effectively distribute agents along the map. As a result, the lowest density is demonstrated exactly by this approach. MARL methods, such as MAMBA and QMIX are also demonstrate low average agents density, as they actually utilize the same observation as Follower does.
Table 9: Average Agents Density by Number of Agents
Algorithm 64 Agents 96 Agents 128 Agents 160 Agents 192 Agents
ASwitcher $0.074$ p m 0.001 $0.111$ p m 0.001 $0.146$ p m 0.001 $0.176$ p m 0.001 $0.203$ p m 0.001
Follower $0.073$ p m 0.000 $0.101$ p m 0.000 $0.126$ p m 0.000 $0.150$ p m 0.000 $0.173$ p m 0.000
IQL $0.080$ p m 0.001 $0.127$ p m 0.002 $0.188$ p m 0.003 $0.257$ p m 0.003 $0.319$ p m 0.002
LSwitcher $0.075$ p m 0.001 $0.114$ p m 0.001 $0.149$ p m 0.001 $0.180$ p m 0.001 $0.208$ p m 0.001
MAMBA $0.073$ p m 0.000 $0.095$ p m 0.000 $0.119$ p m 0.001 $0.146$ p m 0.001 $0.176$ p m 0.001
MATS-LP $0.110$ p m 0.002 $0.176$ p m 0.002 $0.231$ p m 0.003 $0.274$ p m 0.003 $0.306$ p m 0.003
QMIX $0.074$ p m 0.000 $0.100$ p m 0.000 $0.126$ p m 0.000 $0.150$ p m 0.001 $0.175$ p m 0.001
QPLEX $0.078$ p m 0.000 $0.114$ p m 0.001 $0.147$ p m 0.001 $0.176$ p m 0.001 $0.216$ p m 0.002
RHCR $0.088$ p m 0.000 $0.125$ p m 0.001 $0.170$ p m 0.002 $0.242$ p m 0.004 $0.314$ p m 0.004
VDN $0.077$ p m 0.001 $0.109$ p m 0.001 $0.141$ p m 0.001 $0.173$ p m 0.001 $0.204$ p m 0.002
C.6 Pathfinding
Table 10: Pathfinding results.
Algorithm Makespan
ASwitcher $340.56$ p m 79.41
Follower $181.00$ p m 20.95
IQL $1825.95$ p m 144.16
LSwitcher $472.64$ p m 119.23
MAMBA $416.45$ p m 136.01
MATS-LP $179.93$ p m 22.45
QMIX $955.54$ p m 200.68
QPLEX $933.74$ p m 199.18
RHCR $179.82$ p m 20.21
VDN $1733.20$ p m 157.96
Pathfinding metric is tailored to indicate how well the algorithm is able to guide am agent to its goal location. As a result, there is actually no need to run the algorithms on LifeLong instances. Instead, they were run on the same set of instances that were utilized for MAPF approaches. The results of this evaluation are presented in Table 10. Again, the best results were obtained by search-based approach – RHCR. Its implementation was slightly modified to work on MAPF instances, when there is no new goal after reaching the current one. Either optimal or close to optimal paths are able to find Follower and MATS-LP. Followers misses optimal paths due to the integrated technique that changes the edge-costs. MATS-LP adds noise to the root of the search tree that might result in choosing of wrong actions. For the approaches from Switcher family it’s actually almost impossible to find optimal paths as they have no information about global map and operate only based on the local observations.
Appendix D Code examples for POGEMA
{listing}
[…]
POGEMA is an environment that provides a simple scheme for creating MAPF scenarios, specifying the parameters of GridConfig. The main parameters are: on_target (the behavior of an agent on the target, e.g., restart for LifeLong MAPF and nothing for classical MAPF), seed – to preserve the same generation of the map; agent; and their targets for scenario, size – used for cases without custom maps to specify the size of the map, density – the density of obstacles, num_agents – the number of agents, obs_radius – observation radius, collision_system – controls how conflicts are handled in the environment (we used a soft collision system for all of our experiments). The example of creation such instance is presenten in Listing D.
Visualization of the agents is a crucial tool for debugging algorithms, visually comparing them, and presenting the results. Many existing MARL environments lack such tools, or have limited visualization functionality, e.g., requiring running the simulator to provide replays, or offering visualizations only in one format (such as videos). In the POGEMA environment, there are three types of visualization formats. The first one is console rendering, which can be used with the default render methods of the environment; this approach is useful for local or server-side debugging. The preferred second option is SVG animations. An example of generating such a visualization is presented in the listing above. This approach allows displaying the results using any browser. It provides the ability to highlight high-quality static images (e.g., as the images provided in the paper) or to display results on a website (e. g., animations of the POGEMA repository on GitHub). This format ensures high-quality vector graphics. The third option is to render the results to video format, which is useful for presentations and videos.
Appendix E POGEMA Toolbox
The POGEMA Toolbox provides three types of functionality.
The first one is registries to handle custom maps and algorithms. To create a custom map, the user first needs to define it using ASCII symbols or by uploading it from a file, and then register it using the toolbox (see Listing E). The same approach is used to register and create algorithms (see Listing E). In that listing, the registration of a simple algorithm is presented, which must includ two methods: act and reset_states. This approach can also accommodate a set of hyperparameters which the Toolbox handles.
{listing}
[…]
{listing}
[…]
{listing}
[…]
Second, it provides a unified way of conducting distributed testing using Dask  [^11] https://github.com/dask/dask and defined configurations. An example of such a configuration is provided in Listing E. The configuration is split into three main sections; the first one details the parameters of the POGEMA environment used for testing. It also includes iteration over the number of agents, seeds, and names of the map (which were registered beforehand). The unified grid_search tag allows for the examination of any existing parameter of the environment. The second part of the configurations is a list of algorithms to be tested. Each algorithm has its alias (which will be shown in the results) and name, which specifies the family of methods. It also includes a list of hyperparameters common to different approaches, e.g., number of processes, parallel backend, etc., and the specific parameters of the algorithm.
The third functionality and third part of the configuration concern views. This is a form of presenting the results of the algorithms. Working with complex testing often requires custom tools for creating visual materials such as plots and tables. The POGEMA toolbox provides such functionality for MAPF tasks out-of-the-box. The listing provides two examples of such data visualization: a plot and a table, which, based on the configuration, provide aggregations of results and present information in a high-quality form, including confidence intervals. The plots and tables in the paper are prepared using this functionality.
Appendix F Extended Related Work
StarCraft Multi-Agent Challenge — The StarCraft Multi-Agent Challenge (SMAC) is a highly used benchmark in the MARL community. Most MARL papers that propose new algorithms provide evaluations in this environment. The environment offers a large set of possible tasks where a group of units tries to defeat another group of units controlled by a bot (a predefined programmed policy). Such tasks are partially observable and often require simple navigation. However, the benchmark has several drawbacks, such as the need to use the slow simulator of the StarCraft II engine, deterministic tasks, and the lack of an evaluation protocol.
Nevertheless, some of these drawbacks have already been addressed. SMAX [39] provides a hardware-accelerated JAX version of the environment, but it cannot guarantee full compatibility since the StarCraft II engine is proprietary software. SMAC v2 [13] solves the problem of determinism, highlighting this issue in the original SMAC environments. Moreover, an evaluation protocol for the SMAC environment is proposed in [15]. Despite these efforts, it’s hard to say that these tasks require the generalization ability of the agent, since the training and evaluation are conducted on the same scenario.
Multi-agent MuJoCo — In MAMuJoCo, the standard tasks involve agents controlling different sets of joints (or a single joint) within a simulated robot. This set of environments is a natural adaptation of the environment presented in the well-known MuJoCo physics engine [55]. These tasks don’t require high generalization abilities or navigation. In the newer version, MuJoCo provides a hardware-accelerated version, forming the basis for Multi-agent BRAX [39], which enhances performance and efficiency.
Google Research Football — Google Research Football [21] is a multi-agent football simulator that provides a framework for cooperative or competitive multi-agent tasks. Despite the large number of possible scenarios in the football academy and the requirement for simple navigation, the tasks are highly specific to the studied domain. Additionally, the number of possible agents is limited. Moreover, the framework offers low scalability, requiring a heavy engine.
Multi-robot warehouse — The multi-robot warehouse environment RWARE [35] simulates a warehouse with robots delivering requested goods. The environment is highly specific to delivery tasks; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol. The best-performing solution [8] in this environment is trained on only 4 agents. The benchmark is highly related to multi-agent pathfinding tasks; however, it doesn’t provide centralized solution integration, which could serve both as an upper bound for learnable decentralized methods and as a source of expert demonstrations.
Level-Based Foraging — Multi-agent environment LBF [35] simulates food collection by several autonomously navigating agants in a grid world. Each agent is assigned a level. Food is also randomly scattered, each having a level on its own. The collection of food is successful only if the sum of the levels of the agents involved in loading is equal to or higher than the level of the food. The agents are getting rewarded by level of food they collected normalized by their level and overall food level of the episode. The game requires cooperation but also the agents can emerge competitve behavior. The environment is very efficiently designed and very simple to set up; however, it doesn’t support procedurally generated scenarios, thus not requiring generalization abilities or an evaluation protocol.
Flatland — The Flatland environment [30] is designed to address the specific problem of fast, conflict-free train scheduling on a fixed railway map. This environment was created for the Flatland Competition [22]. The overall task is centralized with full observation; however, there is an adaptation to partial observability for RL agents. Unfortunately, during several competitions, despite the presence of stochastic events, centralized solutions [24] from operations research field have outperformed RL solutions by a large margin in both quality and speed. The environment is procedurally generated, which requires high generalization abilities, and the benchmark provides an evaluation protocol. A significant drawback is the extremely slow speed of the environment, which highly restricts large-scale learning.
Overcooked — The Overcooked is a benchmark environment for fully cooperative human-AI task performance, based on the widely popular video game [7]. In the game, agents control chefs tasked to cook some dishes. Due to possible complexity of the cooking process, involving multistep decision-making, it requires emergence of cooperative behaviour between the agents.
Griddly — This is a grid-based game engine [4], allowing to make various and diverse grid-world scenarios. The environment is very performance efficient, being able to make thousands step per second. Moreover, there is test coverage and continuous integration support, allowing open-source development. The engine provides support for different observation setups and maintains state history, making it useful for search based methods.
Multi-player game simulators — Despite the popularity of multi-player games, it’s a challenging problem to develop simulators of the games that could be used for research purposes. One of the most popular adaptations are MineCraft MALMO [19] that allows to utilise MineCraft as a configurable research platform for multi-agent research and model various agents’ interactions. In spite of the game’s flexible functionality, it depends on external runtime, so might be very hard to set up or extremely slow to iterate with. That’s why there are several alternatives that prioritise fast iteration over the environment complexity, like Neural MMO [51] that models a simple MMO RPG with agents with a shared resource pool. On top of that, there are even faster implementation, targeting coordination or cooperation, like Hide-and-Seek [3], modelling competition, or GoBigger [69], focusing on competition between cooperating populations.
Multi-agent Driving Simulators — Autonomous driving is one of the important practical applications of MARL, and Nocturne  [57] is a 2D simulator, written in C++, that focuses on different scenarios of interactions — e.g. intersections, roundabouts etc. The simulator is based on trajectories collected in real life, so allows modelling practical scenarios. This environment has evaluation protocols and supports open-source development with continuous integration and covered by tests. There are also environments, focusing on particular details of driving, for example, Multi-car Racing  [44] that represents racing from bird’s eye view.
Suits of multi-agent environments — These multi-agent environments are designed to be very simple benchmarks for specific tasks. Jumanji  [5] is a set of environments for different multi-agent scenarios connected to combinatorial optimization and control, for example, routing or packing problems. With the purpose for each environment to be focused on the particular task, the overall suit doesn’t test generalization or enable procedural generation. Multi Particle Environments (MPE)  [28] is a communication oriented set of partially observable environments where particle agents are able to interact with fixed landmarks and each other, communicating with each other. SISL  [16] is a set of three dense reward environments was developed to have a simple benchmark for various cooperative scenarios. For environment suits, testing generalization, MeltingPot  [1] comes into place. This set of the environments contains a diverse set of cooperative and general-sum partially observable games and maintains two populations of agents: focal (learning) and visiting (unknown to the environment) to benchmark generalization abilities of MARL algorithms. The set in based on the own game engine and might be extended quite easily.
Real-world Engineering in Practice — Real-world engineering tasks can often be addressed by sophisticated MARL solutions. IMP-MARL [23] provides a platform for evaluating the scalability of cooperative MARL methods responsible for planning inspections and repairs for specific system components, with the goal of minimizing maintenance costs. At the same time, agents must cooperate to minimize the overall risk of system failure. MATE [34] addresses target coverage control challenges in real-world scenarios. It presents an asymmetric cooperative-competitive game featuring two groups of learning agents, cameras and targets, each with opposing goals.
Appendix G Examples of Used Maps
The examples of used maps are presented in Figure 9, showing a diverse list of maps. The map types used in the POGEMA Benchmark include: Maze, with prolonged 1-cell width corridors requiring high-level cooperation; Random, easily generated maps to avoid overfitting with controllable obstacle density; MovingAI-tiles, smaller modified slices of MovingAI maps; Puzzle, small hand-crafted maps with challenging patterns necessitating agent cooperation; Warehouse, used in LifeLong MAPF research, featuring high agent density and throughput challenges; and MovingAI, benchmark maps with varying sizes and structures for single-agent pathfinding.
(a) Maze
(b) Random
(c) MovingAI-tiles
(d) Puzzle
(e) Warehouse
(f) MovingAI
Figure 9: Examples of maps presented in the POGEMA Benchmark. The MovingAI map (on which the pathfinding metric was tested) is shown without grid lines and agents for clarity.
Appendix H MARL Training Setup
For training MARL approaches, such as MAMBA, QMIX, QPLEX, and VDN, we used the default hyperparameters provided in the corresponding repositories [^12] https://github.com/jbr-ai-labs/mamba, and employed the PyMARL2 framework [^13] https://github.com/hijkzzz/pymarl2 to establish MARL baselines. As input, we apply preprocessing from the Follower approach, which is the current state-of-the-art for decentralized LifeLong MAPF. We attempted to add a ResNet encoder, as used in the Follower approach; however, this addition worsened the results, thus we opted for vectorized observation and default MLP architectures. For centralized methods that work with the state of the environment (e.g., QMIX or QPLEX), we utilized the MARL integration of POGEMA, which provides agent positions, targets, and obstacle positions in a format similar to the SMAC environment (providing their coordinates).
Our initial experiments on training this approach with a large number of agents, similar to the Follower model, showed very low results. We adjusted the training maps to be approximately $16\times 16$, which proved to be more effective and populated them with $8$ agents. This setup shows better results. We continued training the approaches until they reached a plateau, which for most algorithms is under 1 million steps.
Appendix I Resources and Statistics
To evaluate all the presented approaches integrated with POGEMA we have used two workstations with equal configuration, that includes 2 NVidia Titan V GPU, AMD Ryzen Threadripper 3970X CPU and 256 GB RAM. The required computation time is heavily depends on the approach by itself.
Table 11: Total time (in hours) required by each of the algorithms to run all MAPF instances on the corresponding datasets.
Random Mazes Warehouse MovingAI-tiles Puzzles MovingAI
DCC 2.11 2.46 11.07 22.70 0.09 0.02
IQL 0.05 0.04 0.13 0.13 0.01 0.01
LaCAM 0.20 0.29 0.24 0.23 0.37 0.01
MAMBA 6.62 6.47 8.36 12.27 2.59 3.40
QMIX 0.04 0.04 0.14 0.13 0.01 0.01
QPLEX 0.05 0.04 0.13 0.13 0.01 0.01
SCRIMP 1.66 2.20 16.54 21.63 0.08 0.21
VDN 0.05 0.04 0.13 0.13 0.01 0.01
Table 12: Total time (in hours) required by each of the algorithms to run all LMAPF instances on the corresponding datasets.
Random Mazes Warehouse MovingAI-tiles Puzzles MovingAI
ASwitcher 1.03 0.47 2.95 1.76 0.31 0.04
Follower 0.48 0.23 0.69 0.77 0.26 0.89
IQL 0.08 0.04 0.26 0.24 0.02 0.01
LSwitcher 6.18 2.61 17.30 10.70 0.81 0.21
MAMBA 13.82 6.69 15.81 11.07 7.83 3.40
MATS-LP 77.31 35.34 163.68 129.78 3.80 0.14
QMIX 0.08 0.04 0.26 0.25 0.02 0.01
QPLEX 0.08 0.04 0.26 0.25 0.02 0.01
RHCR 0.57 0.25 17.04 6.28 0.01 0.01
VDN 0.08 0.04 0.25 0.25 0.02 0.01
The statistics regarding the spent time on solving MAPF and LMAPF instances are presented in Table 11 and Table 12 respectively. Please note, that all these approaches were run in parallel in multiple threads utilizing dask, that significantly reduces the factual spent time.
We used pretrained models for all the hybrid methods, such as Follower, Switcher, MATS-LP, SCRIMP, and DCC, thus, no resources were spent on their training. RHCR and LaCAM are pure search-based planners and do not require any training. MARL methods, such as MAMBA, QPLEX, QMIX, IQL, and VDN, were trained by us. MAMBA was trained for 20 hours on the MAPF instances, resulting in 200K environment steps, and for 6 days on LifeLong MAPF instances, resulting in 50K environment steps, which corresponds to the same amount of GPU hours. For MARL approaches, we trained them for 1 million environment steps, which corresponds to an average of 5 GPU hours for each algorithm.
Appendix J Community Engagement and Framework Enhancements
Our team is committed to maintaining an open and accountable POGEMA framework. Since 2021, we have continuously improved POGEMA, including the addition of the POGEMA Toolbox and the recent introduction of POGEMA Benchmark. We ensure transparency in our operations and encourage the broader AI community to participate. Our framework includes a fast learning environment, problem instance generator, visualization toolkit, and automated benchmarking tools, all guided by a clear evaluation protocol. We have also implemented and evaluated multiple strong baselines that simplify further comparison. We practice rigorous software testing and conduct regular code reviews. We promptly address issues reported on GitHub and welcome any feedback and contributions through GitHub.